### v1

# ILRe: Intermediate Layer Retrieval for Context Compression in Causal Language Models

Abstract
Large Language Models (LLMs) have demonstrated success across many benchmarks. However, they still exhibit limitations in long-context scenarios, primarily due to their short effective context length, quadratic computational complexity, and high memory overhead when processing lengthy inputs. To mitigate these issues, we introduce a novel context compression pipeline, called Intermediate Layer Retrieval (ILRe), which determines one intermediate decoder layer offline, encodes context by streaming chunked prefill only up to that layer, and recalls tokens by the attention scores between the input query and full key cache in that specified layer. In particular, we propose a multi-pooling kernels allocating strategy in the token recalling process to maintain the completeness of semantics. Our approach not only reduces the prefilling complexity from $O(L^{2})$ to $O(L)$, but also achieves performance comparable to or better than the full context in the long context scenarios. Without additional post training or operator development, ILRe can process a single $1M$ tokens request in less than half a minute (speedup $\approx 180\times$) and scores RULER-$1M$ benchmark of $\approx 79.8$ with model Llama-3.1-UltraLong-8B-1M-Instruct on a Huawei Ascend 910B NPU.
1 Introduction
Large Language Models (LLMs) have achieved remarkable success across a wide range of context processing tasks, including document retrieval (Laban et al., 2023), code generation (Gu, 2023), multimodal applications (Sapkota et al., 2025), and agent systems  (Guo et al., 2024; Luo et al., 2025). Especially in the long context scenarios, many works (Liu et al., 2025) are contributed to explore the limits of abilities of LLMs. Meanwhile, mainstream LLMs are increasingly supporting longer context lengths (Yang et al., 2025; Xu et al., 2025; MiniMax et al., 2025).
Method 32K 64K 128K 512K 1M
FullContext 83.2 76.7 70.2 OOM OOM
StreamingLLM 17.7 13.5 OOM OOM OOM
SnapKV 72.4 65.4 OOM OOM OOM
RAG 69.7 67.3 57.9 48.1 48.4
ILRe(Ours) 88.3 84.2 84.0 84.1 79.8
Table 1: RULER benchmark performance of different methods with Llama-3.1-UltraLong-8B-1M-Instruct. Except the FullContext, the token budget is set to be 4K. The best scores are in bold faces.
However, LLMs face significant limitations in long-context applications. First, their effective context lengths are often much shorter than their claimed capacities. For instance, as shown in Table 1, a model claiming to support $1M$ tokens achieves $83.2$ accuracy at $32K$ tokens but drops sharply to $70.2$ at $128K$ in RULER benchmarks (Hsieh et al., 2024). Second, the quadratic computational complexity of attention leads to impractical prefilling times—reaching minutes for long contexts, as demonstrated in Figure 1. Finally, the memory overhead of key-value(KV) caches becomes prohibitive: for example, a $1M$-token cache with Llama-3.1-8B  (Grattafiori et al., 2024) requires $\approx 120GB$ of memory, far exceeding the capacity of one typical GPU/NPU, even though the model already adapts GQA (Ainslie et al., 2023) technique to reduce the KV cache size.
Figure 1: Time To First Token(TTFT) of Intermediate Layer Retrieval (ILRe) with Llama-3.1-8B on a Huawei Ascend 910B NPU. The red text notate the speedup of our approach.
Recently, many works have been contributed to mitigate the above issues. SnapKV (Li et al., 2024b) and Quest (Tang et al., 2024) proposed using the attention scores of an observation window to select salient tokens from the KV cache to reduce the complexity of decoding phase. DuoAttention (Xiao et al., 2025) introduced a streaming chunked prefill to encode incoming tokens within streaming heads in the prefilling phase. However, integrating these techniques into a unified, highly accurate, and efficient pipeline remains an open challenge. Meanwhile, our empirical observations reveal that different decoder layers exhibit varying recall rates of deserved content retrieved by SnapKV-like operation in which we use maximums of attention weights to unify focused content. As illustrated in the “One Kernel” series in Figure 2, some intermediate layers have high recall rates which also exhibit degradation when the length of key increases (same series in Figure 3).
Figure 2: Recall of layers in model Llama-3.1-UltraLong-8B-1M-Instruct over 128K context. The context is prefilled by streaming chunked prefill and the scores are averaged over multiple depths and key lengths.
Figure 3: Recall of different Key Lengths in model Llama-3.1-UltraLong-8B-1M-Instruct over 128K context. The context is prefilled by streaming chunked prefill and the scores are averaged over multiple depths.
Inspired by the works and observations mentioned above, we propose a novel context compression method, Intermediate Layer Retreival(ILRe), as depicted in Figure 4. Our method processes the long full context with partial activated model parameters and streaming chunked prefill to reduce the prefilling complexity. In the layer determined by a simple retrieval task, it collects all the past key states of each chunk and the query states of question so that it can perform a SnapKV-like operation to get full attention weights for query-aware global analysis. Specifically, we propose a strategy that evenly gathers retrieval tokens from multiple combinations of max- and avg-pooling to enhance the capability of retrieving long contents (“Multi Kernels” series in Figure 2 and  3).
Figure 4: An Overview of Intermediate Layer Retreival(ILRe) framework.
We evaluate our approach with various benchmarks, including RULER and LongBench (Bai et al., 2023), to demonstrate its effectiveness. The results show that our approach achieves remarkable performance across all benchmarks over other compression methods.
2 Related Works
The limitations of large language models in long-context scenarios have inspired a range of approaches that aim to reduce memory usage, lower computation complexity, and preserve relevant information across extended sequences. Prior efforts broadly fall into the following two categories: internal memory compression and prompt-level context reduction.
2.1 Internal Memory Compression
Autoregressive LLMs store key–value (KV) pairs during decoding to avoid recomputation, but this cache grows linearly with sequence length, creating a significant memory bottleneck. Several approaches have been proposed to mitigate this issue by selectively retaining or compressing the stored representations. H2O (Zhang et al., 2024) preserves only “heavy hitter” tokens that dominate attention scores, while FastGen (Ge et al., 2024) applies runtime profiling to dynamically choose compression policies. SnapKV (Li et al., 2024c) pools attention scores across prompt tokens to identify and retain informative clusters. Other works identify attention heads specialized for long-range recall and introduce head-specific cache compression (Wu et al., 2024; Fu et al., 2024b; Liu et al., 2024a; Jiang et al., 2024a).
Beyond KV cache optimization, other efforts compress or reuse representations stored in intermediate layers. For example, layer-skipping approaches (Liu et al., 2024b) maintain a smaller subset of intermediate states for recall, balancing compression and semantic completeness. These researches align with other studies that examine the language model internal workflows(Liu et al., 2019), which show that intermediate layers captures increasingly abstract language patterns - from basic grammar to deeper meaning. Such insights motivate designs that exploit the representational richness of intermediate layers while reducing the memory and computation cost of long-context inference.
2.2 Prompt-Level Context Reduction
Prompt compression methods aim to reduce input length before model inference. Semantic compression approaches such as AutoCompressor (Chevalier et al., 2023) and gist-token-based prefix tuning (Mu et al., 2024) learn condensed prompt representations that preserve task-relevant semantics. Meanwhile, token-deletion techniques work at the token level, using metrics such as mutual information (Church & Hanks, 1989) or perplexity (Jiang et al., 2023) to prune less informative tokens.
Beyond these, recent work has proposed reinforcement learning (RL)-based, LLM scoring-based, and LLM annotation-based paradigms for prompt compression. KiS (Laban et al., 2021) addresses the simplification of unsupervised text by balancing fluency, significance, and simplicity through reinforcement learning. Using a k-SCST algorithm, it generates multiple candidate simplifications, evaluates them with a composite reward, and promotes candidates that exceed the mean reward. Similarly, SCRL (Li et al., 2024a) frames sentence compression as a sequence labeling task, fine-tuning a pre-trained transformer with a policy gradient method to classify tokens as essential or non-essential while optimizing for fluency and faithfulness. The selective context (Fu et al., 2024a) represents an LLM scoring-based approach that computes self-information for lexical units using a base causal language model, pruning tokens with low informativeness to yield a concise yet informative prompt. LLMLingua (Jiang et al., 2023) adopts a coarse-to-fine compression pipeline with a budget controller, iterative token-level pruning, and instruction tuning to maintain semantic integrity under high compression ratios. LongLLMLingua (Jiang et al., 2024b) extends this approach for long-context scenarios, incorporating question-aware compression, document reordering to reduce position bias (Yu et al., 2025), and dynamic compression ratios. LLMLingua-2 (Pan et al., 2024) further generalizes the method by introducing a GPT-4-distilled extractive compression dataset and defining compression as a Transformer-based token classification task that uses the full bidirectional context. In addition, a recent study proposes a retrieval framework that uses intermediate representations to obtain external information, which is then incorporated into the prompt to improve the final generation, showing consistent gains in multiple open-source LLMs(Lin et al., 2025).
3 Methodology
In this section, we formally introduce our context compression method, ILRe. In this work, we focus on the decoder-only LLMs.
3.1 Preliminaries
The LLM uses its tokenizer to convert the words to input IDs, which are indices of the embedding metrics $E\in\mathbb{R}^{V\times d}$ of LLMs. Let $\mathbf{x}=(x_{1},x_{2},...,x_{L})$ be the input ids of length $L$. The goal of prompt compression is to generate a subset $\mathbf{\tilde{x}}=\{\tilde{x}_{i}\}_{i=1}^{\tilde{L}}$ of $\mathbf{x}$, where $\tilde{L}<L$ is the allocation budget.
Collect the embeddings to form a input matrix, $X_{1}\in\mathbb{R}^{L\times d}$. This matrix is the initial tokens of the first layer of LLM and then each layer will output a matrix of the same shape as the input of the next layer. For each head $h\in\{1,2,...,H\}$ in layer $l\in\{1,2,...,N_{L}\}$ with head dimension $d_{h}$, we focus on the Query and Key states, which are converted from tokens by three linear transformation matrices $W_{h,l}^{Q}$, $W_{h,l}^{K}$ $\in\mathbb{R}^{d\times d_{h}}$ separately:
$Q_{h,l}=PE(X_{l}W_{h,l}^{Q}),K_{h,l}=PE(X_{l}W_{h,l}^{K})$ (1)
where $PE$ stands for the position encoding. The attention is calculated as:
$A_{h,l}=\text{softmax}(\frac{Q_{h,l}K_{h,l}^{T}}{\sqrt{d_{h}}}\bigodot M_{h,l})$ (2)
where $\{M_{h,l}\}$ are attention masks. In a causal model, $M_{h,l}$ is usually a lower triangular matrix. But with the streaming chunked prefilling (Xiao et al., 2025), $M_{h,l}$ will be a $\Lambda$-like mask as demonstrated in Figure 4.
3.2 Observations of Retrieval Tasks
As demonstrated by SnapKV, the critical components of the KV cache for a specific query can be identified through the attention between the context and the query, which is calculated as:
$A_{h,l}^{q,C}=\text{softmax}\left(\frac{q_{h,l}{K_{h,l}^{C}}^{T}}{\sqrt{d_{h}}}\right)$ (3)
Here, we assume that the batch size is $1$, that is, $A_{h,l}^{q,C}$ is a 3D matrix with head, query, and context dimensions. Let $L_{q}$ denote the size of the query $q$; typically, $L_{q}$ is much smaller than the length $L$ of the context $C$. In the process of context compression, we need to determine which input IDs of context (rather than KV pairs) should be retained for re-inference. To achieve this, we must reduce the attention from 3D to a 1D matrix with only the context dimension left. For that, we only keep the maximums of attention across the head and query dimensions in one specific layer $l_{R}$:
$A^{C}=\max_{h,q}A_{h,l_{R}}^{q,C}$ (4)
Since we calculate the full attention in the layer $l_{R}$ only, we omit the symbol $l_{R}$ for convenience, and, in practice, only the key states will be cached with full length of $L$ in that layer.
SnapKV also noted that using average pooling preserves semantic features better than simply retaining only those with the highest attention weights. Here, we extend this technique by incorporating max pooling, multiple kernel sizes and budget allocation. This procedure is detailed in Algorithm 1. SnapKV constitutes a special case of this algorithm, featuring one max pooling kernel of size $1$ and one average pooling kernel. By default, we use $3$ max pooling kernels with sizes 2,4,8{2,4,8} and $16$ average pooling kernels with sizes from $1$ to $16$. This results in 48 combinations of max and average pooling. We evenly distribute the token budget across all combinations. If a token has already been allocated by previous combinations, the next-highest-ranked candidate is selected instead, continuing until the target budget is met.
Algorithm 1 Context Allocation
Input: attention weights $A^{C}$ without sink part, max pooling kernel sizes $\{m_{i}\}_{i=1}^{N_{M}}$, average pooling kernel sizes $\{n_{j}\}_{j=1}^{N_{A}}$, total budget $B$, input IDs $C$;
Output: compressed context $\tilde{C}$
 Init a set of allocated indices II with indices of sink part
 Calculate the budget of each combination $B_{c}=\frac{B}{N_{M}N_{A}}$
for $i$ $\leftarrow$ 11 to $N_{M}$ do
  Calculate the k of top in this max kernel $K_{m}=\frac{B}{m_{i}}+1$
  Calculate $\tilde{A}^{C}_{m_{i}}=\text{MaxPooling}(A^{C})$ with $size=m_{i}$, $stride=m_{i}$
  for $j$ $\leftarrow$ 11 to $N_{A}$ do
   Calculate $\tilde{A}^{C}_{m_{i},n_{j}}=\text{AvgPooling}(\tilde{A}^{C}_{m_{i}})$ with $size=n_{j}$, $stride=1$
   Select top $K_{m}$ indices from $\tilde{A}^{C}_{m_{i},n_{j}}$ as $\tilde{I}_{m_{i},n_{j}}$
   Map $\tilde{I}_{m_{i},n_{j}}$ back to indices of $C$ as $I_{m_{i},n_{j}}$
   Let $S_{I}=0$
   while $S_{I}<B_{c}$ do
    Get next index $I^{k}_{m_{i},n_{j}}\leftarrow I_{m_{i},n_{j}}$
    if $I^{k}_{m_{i},n_{j}}\not\in I$ then
     Update $I=I\cup I^{k}_{m_{i},n_{j}}$
     Update $S_{I}=S_{I}+1$
    end if
   end while
  end for
end for
 Sort II in ascending order
 Gather input IDs from $C$ by indices II as new context $\tilde{C}$ Return compressed context $\tilde{C}$
In the simple retrieval task defined in Appendix A.1, we calculate the recall rate of each layer across varying key depths and lengths, as presented in Figure 2 and Figure 3. We observe that certain layers equipped with only a single average kernel can achieve a $100\%$ recall rate in short-key tasks but decline sharply as the key size increases. In contrast, our multi-kernel approach maintains consistent performance across different key sizes. Furthermore, frontier layers exhibit higher recall rates than later layers, resulting in the optimal $l_{R}$ typically being smaller than half of $N_{L}$ and thus yielding much lower computational complexity. We define the optimal layer $l_{R}$ as the layer with the smallest index value among those exhibiting the highest recall rates. For example, in the model Llama-3.1-Ultralong-8B-1M-Instruct (Xu et al., 2025), we identify the optimal $l_{R}=3$ versus the total number of layers $N_{L}=32$. Meanwhile, we count the frequency of occurrence of maxima arguments of attention heads in Figure 5, which demonstrates that attention heads are not fixed in their behavior.
Figure 5: Occurrence Ratios of max head indices in one $128K$ Retrieval Task of model Llama-3.1-Ultralong-8B-1M-Instruct.
3.3 Intermediate Layer Retrieval
The main schema of Intermediate Layer Retrieval(ILRe) is present in Figure 4. For end-to-end inference:
  • Determine $l_{R}$ offline as discussed above.
  • Load LLM parameters up to the layer $l_{R}$. Only these parameters will be used for the context compression.
  • Prepare a long context and query part. Usually, the query part consists of a question and answer prefix.
  • In layers $\{1,2,...,l_{R}-1\}$, apply the streaming chunked prefill. Only cache the sink and sliding part of the key states and value states with sizes $S$ and $W$ respectively.
  • In the retrieval layer $l_{R}$, calculate the key states for the context and the query states for the query part, and then skip the rest of the calculations. Pre-allocate memory with context length $L$ for key states caching $K^{C}_{h,l_{R}}$ and fill the cache chunk by chunk.
  • Prefill the query part to get the query states $q_{h,l_{R}}$.
  • Calculate $A^{C}$ with the cached key states and the query states.
  • Apply algorithm 1 to select input IDs.
  • Concatenate the selected input IDs with those of the query part to form a new input for standard inference. This component can be offloaded to other highly optimized inference frameworks, such as vLLM (Kwon et al., 2023), SGLang (Zheng et al., 2024), and others.
3.4 Complexity Analysis
We only discuss the attention complexity of context compression, as the standard inference of a compressed prompt is constrained by a fixed budget $B$.
Method Computation Memory
FlashAttention $O(N_{L}L^{2})$ $O(2N_{L}L)$
ILRe(Ours) $O((W^{{}^{\prime}}l_{R}^{{}^{\prime}}+L_{q})L)$ $O(2W^{{}^{\prime}}l_{R}^{{}^{\prime}}+L)$
Table 2: Complexity of Computation and Memory footprint. For convenience, we set $W^{{}^{\prime}}=S+W$ and $l_{R}^{{}^{\prime}}=l_{R}-1$. The factor 22 is for the key and value states.
The complexity of our approach is listed in Table 2. Compared to the standard FlashAttention (Dao, 2023), ILRe only calculates attentions in a streaming chunk fashion in $l_{R}-1$ layers and in an observation window manner in the specific layer $l_{R}$. As $S,W,L_{q}\ll L$ in the long context scenarios, the computation complexity of ILRe approximates $O(L)$; notably, the memory footprint is roughly $\frac{1}{2N_{L}}$ of that of the full context.
4 Experiments
Single Doc. QA Multi Doc. QA Summarization Fewshot Learning Synthetic Code
$B$ Method
NarrativeQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PR-en
Lcc
RB-P
Avg.
FullContext 23.5 45.4 54.3 44.8 41.6 24.1 34.4 23.6 27.4 69.5 91.2 44.4 6.6 76.0 63.2 54.9 45.3
2K SnapKV 30.8 42.8 52.3 53.4 44.5 29.3 30.3 24.8 27.0 2.0 91.6 8.3 4.5 99.5 18.6 15.0 35.9
StreamingLLM 22.9 34.3 37.1 43.9 34.4 18.0 28.1 22.2 26.8 0.0 72.8 8.4 2.4 95.0 14.6 12.9 29.6
RAG 18.6 41.8 48.4 45.5 40.8 24.7 30.9 23.1 26.9 67.0 91.0 40.7 2.0 67.0 57.8 49.1 42.2
ILRe w/ $l_{R}=3$ 22.0 42.0 54.6 53.1 34.7 26.8 30.7 22.3 27.1 71.0 91.2 37.9 4.8 72.5 63.9 55.6 44.4
ILRe w/ $l_{R}=6$ 20.4 43.8 55.2 57.5 41.6 27.3 31.1 24.0 27.0 69.0 90.6 43.2 3.1 76.5 59.6 49.8 45.0
ILRe w/ $l_{R}=10$ 22.0 43.5 50.9 50.7 39.2 26.7 31.4 20.6 27.2 68.0 89.6 37.8 3.9 83.5 63.6 56.3 44.7
ILRe w/ $l_{R}=11$ 26.0 45.4 53.6 55.3 40.2 29.2 31.5 22.4 27.1 73.5 89.9 38.2 4.6 83.0 64.3 56.9 46.3
[…]
1K SnapKV 29.4 40.9 51.8 53.1 45.7 29.2 27.6 24.0 26.0 0.5 91.4 11.3 4.0 99.5 18.3 16.1 35.6
StreamingLLM 22.4 27.0 35.6 42.2 26.1 15.4 25.6 21.6 25.9 0.0 70.5 9.4 2.8 97.0 11.6 15.2 28.0
RAG 20.5 35.4 45.4 46.7 40.6 22.9 28.0 22.0 25.9 62.0 87.4 38.0 2.1 46.0 53.3 48.1 39.0
ILRe w/ $l_{R}=3$ 18.1 33.2 51.5 49.2 34.8 21.6 28.6 20.1 26.2 66.0 85.7 37.6 2.5 36.5 63.6 54.4 39.4
ILRe w/ $l_{R}=6$ 19.0 40.2 53.7 56.8 36.0 23.4 29.0 22.8 25.8 66.5 89.0 41.6 2.9 45.5 54.7 50.1 41.1
ILRe w/ $l_{R}=10$ 19.3 35.9 43.1 43.1 29.9 16.7 29.6 17.3 26.4 59.5 79.4 36.6 4.6 46.5 61.2 53.6 37.7
ILRe w/ $l_{R}=11$ 23.4 44.2 50.2 50.9 37.0 23.7 29.2 19.5 26.0 70.0 85.2 36.4 2.6 56.0 62.3 54.3 41.9
[…]
Table 3: Longbench results of model Llama-3.1-8B-Instruct. The best and second best scores are boldfaced and underlined respectively.
In this section, we empirically evaluate the effectiveness of our approach, ILRe.
Baseline Methods. For comparison, except the original prompt with full context, we select other methods similar to or used in the components of our approaches: StreamingLLM (Xiao et al., 2023), SnapKV (Li et al., 2024b), and Retrieval-Augmented Generation (RAG) (Gao et al., 2024; Gupta et al., 2024) with Qwen3-Embedding-0.6B (Zhang et al., 2025) as the backbone.
Benchmarks and LLMs. In this work, we utilize two benchmarks: LongBench and RULER. For RULER, we set the token budget to $4K$. Given that LongBench contexts are relatively short (with an average length of fewer than $16K$ tokens), we only test the Llama-3.1-8B-Instruct model (Grattafiori et al., 2024)—which claims a $128K$ context capacity—setting the token budget to $1K$ and $2K$. We also test an additional model family: Qwen2.5-7B-Instruct (Qwen et al., 2025). Similar to Figure 2 and  3, the results of Llama-3.1-8B-Instruct and Qwen2.5-7B-Instruct in the Retrieval Task are present in Figure 6 and  7 respectively.
Figure 6: Results of model Llama-3.1-8B-Instruct with $128K$ context in Retrieval Trask.
Figure 7: Results of model Qwen2.5-7B-Instruct with $32K$ context in Retrieval Trask.
Default Settings of ILRe. By default, we set $S=4$, $W=512$, and the chunk size of prefilling is 10241024(the first chunk with additional $S$). For the pooling, we use $3$ max pooling kernels with sizes $2,4,8$ and $16$ average pooling kernels with sizes from $1$ to $16$ as mentioned in Section 3.2.
4.1 RULER
The RULER results are presented in Table 1. Our approach consistently outperforms other methods, including the full context with original prompt. Furthermore, performance degradation exhibits a much weaker dependence on context length compared to other methods.
4.2 LongBench
We evaluate our method on the LongBench benchmark, which measures long-context understanding across a diverse set of tasks, including question answering, summarization, fewshot learning, synthetic tasks, and coding. Table 3 reports the results under token budget of 2K and 1K.
We highlight two key observations. First, ILRe achieves improvements over the full context at 2K token budget. ILRe with full KV retention attains the best overall performance, outperforming SnapKV and StreamingLLM in most tasks, with particularly strong gains in retrieval-heavy tasks such as QA and summarization. Even with various layers (e.g. $l_{R}=3,6,10,11$), ILRe remains competitive, demonstrating that intermediate layer retrieval can recover useful long-context information with fewer stored tokens. Second, at 1K context length, ILRe continues to scale favorably, showing robustness under tighter memory budgets.
4.3 Ablations
In this part, we will change the settings or modify the components of the standard ILRe to prove its robustness.
Settings 32K 64K 128K
Streaming Chunked Prefill 88.3 84.2 84.0
FullKV Prefill 87.2 83.5 84.0
Table 4: RULER benchmark performance of different prefilling methods with Llama-3.1-UltraLong-8B-1M-Instruct.
Prefilling Method. To prove the stability of the streaming chunked prefill in our framework, we compare it with using the full KV in layers $\{1,2,...,l_{R}-1\}$ by setting $W=\infty$. As demonstrated in Table 4, the results of RULER benchmark of the streaming chunked prefill are comparable to those of full KV. The results with “FullKV” in Table 3 further reinforce this finding.
Settings 128K 512K
NS2 NS3 Avg. NS2 NS3 Avg.
Default 100.0 100.0 84.0 100.0 100.0 84.1
Less-Max-Kernels 100.0 99.2 82.4 100.0 99.4 82.8
Less-Avg-Kernels 100.0 100.0 82.5 100.0 100.0 83.1
One-Max-Avg 100.0 0.0 65.6 100.0 1.2 66.6
Table 5: RULER benchmark performance of different pooling settings with Llama-3.1-UltraLong-8B-1M-Instruct. The best of average scores are in bold faces.
Multi Kernels. We define three different configurations of the pooling method from the default setups:
  • Less-Max-Kernels uses only one max pooling with kernel size $4$.
  • Less-Avg-Kernels reduces the number of average pooling to nine with sizes ranging from $1$ to $9$.
  • One-Max-Avg only has one max pooling and one average pooling with size $4$ and $5$ respectively.
These tasks are evaluated at $128K$ and $512K$ in the RULER benchmarks present in Table 5. We include two subtasks, NIAH-Single-Key-2 (NS2) and NIAH-Single-Key-3 (NS3), to demonstrate that multi-kernels are effective for long-content retrieval. This is evident as the One-Max-Avg setup exhibits an abrupt performance drop in the longer-key task NS3.
8K 16K 32K
FullContext 93.1 91.4 84.7
ILRe with $l_{R}=6$ 87.2 85.1 69.6
ILRe with $l_{R}=5$ 93.6 91.9 87.9
Table 6: RULER benchmark performance of ILRe with Qwen2.5-7B-Instruct. The best scores are in bold faces.
Other Models. For the model Qwen2.5-7B-Instruct, we identify the optimal $l_{R}=5$ and limit testing to $32K$ in Table 6, consistent with its claimed capabilities. This result demonstrates that our approach is valid across other model families.
4.4 Efficiency
In Figure 1, we measure the Time To First Token (TTFT) of our approach with a batch size of one on a single Huawei Ascend 910B NPU, averaging results across multiple trials. Due to hardware limitations, we only measured the full-context baseline up to $256K$ tokens, and extended the results to $1M$ tokens using a quadratic fit. Given the large discrepancy in TTFT between our approach and the full-context baseline, we use the log-scale y-axis in the figure. Our method demonstrates a significant speedup, achieving up to a hundredfold improvement in long-context inference. Additionally, the linear fit of ILRe yields an $R^{2}$ value of $0.994$, which strongly aligns with our complexity analysis conclusion of $O(L)$ in Section 3.4.
5 Conclusion
In this study, we propose a training-free context compression framework, ILRe. Built upon standard causal LLMs, it is highly efficient and effective in context prefilling and extraction. By integrating streaming chunked prefill, an observation window, multi pooling kernels and an early exiting layer, ILRe reduces the computational overhead of self-attention, delivering up to a hundredfold speedup in end-to-end inference while maintaining remarkable accuracy. Our experimental evaluations demonstrate that ILRe holds significant promise for advancing long-context LLM inference.
Appendix A Appendix
A.1 Details of Retrieval Task
In the retrieval task for determining $l_{R}$, we embed a number of numerical passkey phrases with varying lengths in random positions within a long textual context. The budget of the retrieval task is set to $1024$. We evaluate performance using exact match accuracy, where a prediction is correct only if the entire passkey phrase is retrieved without error. Following the Needle-in-a-Haystack setup (Kamradt, 2023), the evaluation spans multiple experimental setups that vary in key depth, key length, and retrieval layer/kernel configuration.
[…]